OpenStudy (anonymous):

Evaluate the summation of 25 times 0.3 to the n plus 1 power, from n equals 2 to 10..

OpenStudy (anonymous):

does anyone know this.. help

OpenStudy (science0229):

So, you can write this question as \[\sum_{n=2}^{10}25*0.3^n*0.3=\sum_{n=2}^{10}7.5*0.3^n\]

OpenStudy (science0229):

We can write that as \[7.5*0.3^2+7.5*0.3^3+...+7.5*0.3^10\], right?

OpenStudy (science0229):

Correction: The last term is supposed to be 7.5*0.3^10

OpenStudy (anonymous):

no this is not right

OpenStudy (science0229):

Why not?

OpenStudy (anonymous):

Evaluate the summation of negative 2 n minus 3, from n equals 2 to 10. then how do u do this in sigma notation

OpenStudy (science0229):

\[\sum_{n=2}^{10}(-2n-3)\]

OpenStudy (anonymous):

yes!!!!!! thankyou how do you solve this

OpenStudy (science0229):

There is a sigma property.

OpenStudy (science0229):

\[\sum_{x=a}^{b}(nx+m)=n \sum_{x=a}^{b}x+\sum_{x=a}^{b}m\]

OpenStudy (science0229):

a,b,n,m are all constants

OpenStudy (anonymous):

okay how do u solve the one i gave you... my answer choices are -270 -135 135 270

OpenStudy (science0229):

You mea the second one?

OpenStudy (anonymous):

yes that one

OpenStudy (science0229):

First, use the sigma property I gave you to get this

OpenStudy (science0229):

\[-2\sum_{n=2}^{10}n-\sum_{n=2}^{10}3\]

OpenStudy (anonymous):

can you show me please

OpenStudy (science0229):

On what part do you need help on?

OpenStudy (anonymous):

i got 135 but i dont think its right

OpenStudy (science0229):

First, do you know how to calculate \[\sum_{n=2}^{10}n\]

OpenStudy (anonymous):

no i really dont. i just learned this

OpenStudy (science0229):

Ok.

OpenStudy (science0229):

If you expand that, you get 2+3+4+5+6+7+8+9+10

OpenStudy (anonymous):

okay i got that idk what to do afterwards

OpenStudy (science0229):

2+3+4+5+6+7+8+9+10=A 10+9+8+7+6+5+4+3+2=A 12+12+12+12+12+12+12+12+12=2A 12*9=2A 54=A

OpenStudy (science0229):

So, \[\sum_{n=2}^{10}n=54\]

OpenStudy (anonymous):

im confussed bc thats not the answer to my problem :(

OpenStudy (science0229):

Going back to what I previously said, we said that \[\sum_{n=2}^{10}(-2n-3)=-2\sum_{n=2}^{10}n-\sum_{n=2}^{10}3\]

OpenStudy (science0229):

We found that \[\sum_{n=2}^{10}n=54\]

OpenStudy (anonymous):

Evaluate the summation of 3 n plus 2, from n equals 1 to 14.. 39 49 340 343

OpenStudy (science0229):

So, the 1st equation becomes \[-2(54)-\sum_{n=2}^{10}3=-108-\sum_{n=2}^{10}3\]

OpenStudy (science0229):

\[\sum_{n=2}^{10}3=3*9=27\]

OpenStudy (science0229):

So, the 2nd equation becomes -108-27, which is -135

OpenStudy (science0229):

So,\[\sum_{n=2}^{10}(-2n-3)=-135\]

OpenStudy (anonymous):

okay i got that !!!!!

OpenStudy (anonymous):

Evaluate the summation of 3 n plus 2, from n equals 1 to 14. how you do this one

OpenStudy (science0229):

Try it on your own. I'll help if you're stuck. You can show work on here.

OpenStudy (anonymous):

okay so do i do the 1+2+3+4 thing

OpenStudy (anonymous):

im not good at this

OpenStudy (science0229):

After that problem, I'll tell you a shortcut to these kind of problems.

OpenStudy (anonymous):

okay thankyou. im sorry to give you all this trouble

OpenStudy (anonymous):

so do i do the 1+2+3 thing

OpenStudy (science0229):

Ok.

OpenStudy (anonymous):

what do i do now?

OpenStudy (science0229):

What did you do so far?